### v1

# Theory of intervalley-coherent AFM order and topological superconductivity in tWSe2

Abstract
The recent observation of superconductivity in the vicinity of insulating or Fermi surface reconstructed metallic states has established twisted bilayers of WSe2 as an exciting platform to study the interplay of strong electron-electron interactions, broken symmetries and topology. In this work, we study the emergence of electronic ordering in twisted WSe2 driven by gate-screened Coulomb interactions. Our first-principles treatment begins by constructing moiré Wannier orbitals that faithfully capture the bandstructure and topology of the system and project the gate-screened Coulomb interaction onto them. Using unbiased functional renormalization group calculations, we find an interplay between intervalley-coherent antiferromagnetic order and chiral, mixed-parity $d/p$-wave superconductivity for carrier concentrations near the displacement field-tunable van-Hove singularity. Our microscopic approach establishes incommensurate intervalley-coherent antiferromagnetic spin fluctuations as the dominant electronic mechanism driving the formation of superconductivity in $\theta=5.08^{\circ}$ twisted WSe2 and demonstrates that nesting properties of the Fermi surface sheets near the higher-order van-Hove point cause an asymmetric density dependence of the spin ordering as the density is varied across the van-Hove line, in good agreement with experimental observations. We show how the region of superconducting and magnetic order evolves within the two-dimensional phase space of displacement field and electronic density as twist angle is varied between $4^{\circ}\dots 5^{\circ}$.
I Introduction
Transition metal dichalcogenide (TMD) heterostructures have arisen as fertile and highly versatile platforms for exploring the interplay between topology and strong electronic correlations Kennes et al. (2021), hosting correlated insulators Wang et al. (2020); Xu et al. (2020); Huang et al. (2021), spin-ordered metals Ghiotto et al. (2024); Anderson et al. (2023); Crépel and Fu (2023), rich multi-orbital physics Zhang et al. (2021); Ryee and Wehling (2023), Feschbach resonances Slagle and Fu (2020); Crépel and Fu (2021); Schwartz et al. (2021); Kuhlenkamp et al. (2022); Crépel et al. (2023); Wagner et al. (2023); Lange et al. (2024), quantum critical behavior Ghiotto et al. (2021); Li et al. (2021), excitonic insulators Tran et al. (2019); Ma et al. (2021); Gu et al. (2022); Xiong et al. (2023), topological Kondo physics Dalal and Ruhman (2021); Guerci et al. (2023); Zhao et al. (2023a); Guerci et al. (2024a); Zhao et al. (2023b); Xie et al. (2024a), quantum anomalous and quantum spin-Hall insulators Tao et al. (2024); Kang et al. (2024a), ferromagnetic fractional quantum anomalous phases Zeng et al. (2023); Cai et al. (2023); Xu et al. (2023); Park et al. (2023), or gapped spin liquids Kang et al. (2024b); Crépel and Millis (2024a). Additionally, early experiments on $5.0^{\circ}$ twisted WSe2 hinted at the existence of superconducting phases in moiré TMDs Wang et al. (2020); a finding recently confirmed by sub-Kelvin transport measurements on $3.65^{\circ}$ Xia et al. (2024) and $5.1^{\circ}$ Guo et al. (2024) twisted samples. The discovery of superconductivity in these highly tunable moiré TMDs has sparked many theoretical studies Schrade and Fu (2021); Wietek et al. (2022); Klebl et al. (2023); Kim et al. (2024); Zhu et al. (2024); Christos et al. (2024); Guerci et al. (2024b); Tuo et al. (2024); Chubukov and Varma (2024); Xie et al. (2024b); Qin et al. (2024); Chubukov and Varma (2024) for the new insights it promises to provide on the origin and properties of strongly correlated topological superconductivity, including exciting prospects for trapping non-abelian anyons at clean, gate-defined interfaces Mong et al. (2014); Vaezi (2014); Barkeshli (2016); Crépel and Regnault (2024).
Twisted WSe2 (tWSe2) consists of two WSe2 layers stacked at a relative twist angle $\theta$ which are encapsulated in a structure including top and bottom gates. The twist, in combination with the electrostatic potential from the local arrangement of WSe2 bilayers and the interlayer electronic hybridization, produces a large unit cell with two dimensional “moiré” bands Wu et al. (2019); Devakul et al. (2021). The average of the potentials on the top and bottom gates tunes the electronic density over wide ranges. The difference in voltage between the top and bottom gates introduces an interlayer potential difference, or “displacement field”, which shifts electronic wave functions and charge density between the layers, modifying the electronic band structure. At a fixed twist angle, the system’s behavior can be explored in a phase diagram defined by electronic density and displacement field. As the density varies at a fixed displacement field, the Fermi surface can pass through a van-Hove singularity (VHS), where the density of states diverges. Of particular relevance for this study is that both the density at which a van-Hove singularity occurs and the strength of the resulting divergence in the density of states depend on twist angle and are tunable over wide ranges by varying the displacement field.
In the $5.1^{\circ}$ tWSe2 device Guo et al. (2024), the superconducting state is observed only for a small range of displacement fields near the hole density $n/n_{0}$ at which the Fermi surface passes through the van-Hove singularity, and is adjacent to a metallic region characterized by an increase in longitudinal resistance and a very rapid non-monotonic doping dependence of the Hall resistivity. This unusual metallic region has been interpreted Guo et al. (2024) as arising from the emergence of antiferromagnetic (AFM) spin ordering that partially gaps out the Fermi surface. The densities at which superconductivity and antiferromagnetism occur are not close to one hole per moiré unit cell and the experimentally important displacement fields are significant, on the order of half of the field required to fully layer-polarize the doped holes. In the $3.65^{\circ}$ device Xia et al. (2024), superconductivity is found only for carrier concentrations near $n/n_{0}=1$ per moiré unit cell as well as small displacement fields, with the superconductivity giving way to an insulating phase and then to an apparently conventional metallic phase as the displacement field is increased. In each device, the superconducting and AFM or insulating states are found only in a relatively small region of the experimental phase space defined by carrier concentration and interlayer (displacement field) potential difference.
The association of superconductivity with proximity to the non-integer carrier concentration at which the Fermi surface contains the VHS, and the proximity to an apparent metallic antiferromagnetic phase suggests a weak-to-moderate coupling scenario in the $5.1^{\circ}$ device, where the effects of electron-electron interactions are amplified by the density of states divergence associated with the van-Hove point. The association of superconductivity at particular values of the displacement field may then be related to proximity to a higher order van-Hove singularity (HOVHS) with a stronger density of states divergence appearing as the band theory is tuned Bi and Fu (2021). Recent literature has presented various versions of this scenario. Interactions considered include spin-fluctuations Klebl et al. (2023); Schrade and Fu (2021); Chubukov and Varma (2024), over-screened Coulomb interactions Guerci et al. (2024b) and phenomenological attraction Zhu et al. (2024); Tuo et al. (2024). While these treatments provide important insights, they are based on particular ansätze and come to different conclusions regarding the symmetry of the superconducting state.
Moreover, theory should take into account the non-trivial topological properties Devakul et al. (2021); Crépel and Millis (2024b); Zhang et al. (2024a) of the low-lying hole bands. In addition to potentially endowing the superconducting state with interesting properties Guerci et al. (2024b), they obstruct an effective single-orbital description of the topmost moiré valence band and imply that interactions projected onto the top-most band generically have a long-ranged structure. A minimal faithful low-energy model of tWSe2 comprises three orbitals per valley that are centered on the triangular and honeycomb sites of the moiré superlattice Crépel and Millis (2024b); Qiu et al. (2023). Such a model accurately accounts for the topological entanglement of the topmost valence bands with Bloch states far below the Fermi energy and for the subtle aspects of the fermiology and its dependence on the displacement field. In this multi-orbital Wannier basis, the gate-screened Coulomb interactions relevant for the experimental setup can be cast into fairly local interactions, which ensures the applicability of established many-body techniques beyond the mean-field level Tuo et al. (2024); Guerci et al. (2024b), including the functional renormalization group Salmhofer and Honerkamp (2001); Metzner et al. (2012); Platt et al. (2013); Dupuis et al. (2021); Beyer et al. (2022) (FRG), dynamical mean-field theory Georges et al. (1996), or tensor network based approaches Schollwöck (2011); Orús (2014). A similar situation arises in twisted bilayers of graphene, where a multi-band “heavy fermion” description is argued to best represent the physics Carr et al. (2019a); Song and Bernevig (2022).
In this work, we go beyond initial phenomenological treatments of correlated phases in tWSe2 and provide a full-fledged analysis of correlated states in the weak-to-moderate interaction regime based on physically realistic repulsive gate-screened Coulomb interactions projected onto a multiband model based on exponentially localized Wannier functions that captures the topology and the intricacies of the band structure. Using unbiased functional renormalization group techniques, we map out the interacting phase diagram of tWSe2 in the twist angle range $\theta=4^{\circ}\dots 5^{\circ}$ from first principles. We demonstrate that intervalley-coherent anti-ferromagnetic (IVC-AFM) order and extended regions of chiral $d/p$-wave superconductivity coexist along the gate-tunable van-Hove line. The unbiased characterization of particle-hole and particle-particle instabilities in conjuction with the subtle fermiology near the van-Hove line enabled by our choice of methods permits us to show that the electronic driving mechanism for superconductivity in tWSe2 is primarily facilitated by spin fluctuations around regions of incommensurate IVC-AFM order below the higher order van-Hove singularity, where the absence of perfect nesting strengthens particle-particle fluctuations. Our microscopic approach can explain several features observed in recent experimental measurements, including the unusual density dependence of the particle-hole instabilities with respect to the van-Hove line. We predict a successive shift of superconducting regions to lower values of the displacement field and towards $n/n_{0}\sim 1$ as the twist angle is decreased and briefly discuss the connection of our results to observations on the $\theta=3.65^{\circ}$ sample, where superconductivity is observed for $n/n_{0}\sim 1$ and for small displacement fields.
Figure 1: Fermiology and Wannierization of $\theta=5.08^{\circ}$ tWSe2 for an external displacement field $E_{z}=20\,\mathrm{meV}$. (a) Continuum bandstructure in valley $\nu=+1$ (black) and spectral overlap of its Bloch states with a three-orbital model comprising Wannier orbitals centered at the triangular ($T$, $s$-orbital) and honeycomb ($H_{1}$, $p_{+}$-orbital)/($H_{2}$, $p_{+}$-orbital) sites of the moiré superlattice (tri-colored map). The bandstructure of the effective Wannier model truncated after 4.5 moiré lattice vectors is indicated by the green dashed line. (b) Amplitude $|\psi_{l}\bm{(}r)|^{2}$ of the Wannier functions ($T,H_{1},H_{2}$) on the two layers ($l=0,1$) that capture the spectral weight of the continuum model. (c) Density of states as function of holes per moiré unit cell $n/n_{0}$ and displacement field $E_{z}$. The upper panel shows snapshots of the Fermi surface along the van-Hove line and its orbital polarization using the same colormap as in (a) for valley $\nu=+1$. The Fermi surface of the opposite valley $\nu=-1$ is shown in gray. The green dashed line indicates the mini-Brillouin zone of tWSe2. The blue circle indicates the position of the higher-order VHS, where three van-Hove points merge at $K^{\nu}$ as demonstrated in the insets. The three van-Hove points are moved towards $\Gamma$ ($M$) for increasing (decreasing) value of the external displacement field along the van-Hove line as indicated by the blue triangle (square).
II Model and method
II.1 Multi-orbital Wannier model
The non-interacting electronic structure of hole-doped tWSe2 is accurately described by a continuum model in which the spin-valley locked holes of both monolayers are treated within the effective mass approximation, and are coupled by moiré potentials that slowly vary on the atomic scale Wu et al. (2019). Due to the emergent spin-valley locking in monolayer WSe2, the low-energy states can be assigned a well-defined valley index $\nu=\pm 1$ and states from opposite spin-valley sectors are related by virtue of time-reversal symmetry $\mathcal{T}$. The effects of out-of-plane electric fields are captured by a layer-dependent potential energy ±$\pm E_{z}$. For twist angles $\theta\gtrsim 3^{\circ}$, in-plane lattice relaxation is weak, and the associated moiré fields are characterized by only a few harmonics Carr et al. (2018); Zhang et al. (2024b); Jia et al. (2024); Thompson et al. (2024); Liu et al. (2024). Here, we use the continuum model from Ref. Devakul et al. (2021), whose coefficients have been fitted to large-scale ab-initio calculations, see Supplementary Materials (SM) SM for details.
As most many-body methods require a symmetry-preserving lattice description of the system under scrutiny, we seek to construct a Wannier representation of the topmost valence band states in each valley based on the three-orbital model put forward in Refs. Crépel and Millis (2024b); Qiu et al. (2023). This model is based on generic symmetry arguments and comprises one triangular lattice site $T$ (at Wyckoff position 1a of the triangular superlattice), which has weight on both layers, as well as two honeycomb lattice sites $H_{1,2}$ (at Wyckoff positions 2a,b) with dominant weight on either one of the layers. At a twist angle of $\theta=5.08^{\circ}$, however, the Chern number sequence of the first three valence bands of tWSe2 in valley $\nu=\pm 1$ is $(\nu,\nu,\nu)$ for the parameters of Ref. Devakul et al. (2021), prohibiting an exponential localization of Wannier functions centered at aforementioned Wyckoff positions constructed from the three topmost bands alone. We resolve this issue by employing a single-shot Wannierization scheme Marzari et al. (2012); Carr et al. (2019b); Fischer et al. (2024) in which Bloch states from the entire valence band manifold are considered for the projection and a proper subspace selection is instead achieved by weighting Bloch states according to their energetic hierarchy and their weight around the desired orbital center. Compared to the layer-polarization scheme presented in Ref. Crépel and Millis (2024b), the energetic subspace selection ensures to construct exponentially localized Wannier functions irrespective of the Chern number of the third moiré band, which sensitively depends on model parameters due to the large energy overlap of this band with remote energy bands. Figure 1 (a) shows the band structure of the Wannier model (dashed cyan lines) compared to the continuum model (black lines) at a displacement field $E_{z}=20\,\mathrm{meV}$. By adding the overlap of each Wannier function with the continuum model’s Bloch states as color, we demonstrate how states from the fourth and fifth bands are essential for a three-orbital description with an overall Chern number sequence $(\nu,\nu,-2\nu)$. The residual spectral weight in remote energy bands is required to render the entire Wannier band manifold topologically trivial. In spite of this, the topmost band of the continuum model is perfectly reproduced in the Wannier model as shown in Fig. 1 (a).
In accordance with previous studies, the topmost band has significant weight in the $T$-orbital Wang et al. (2020); Pan et al. (2020); Zang et al. (2022); Kiese et al. (2022); Klebl et al. (2023); Ryee and Wehling (2023); Zegrodnik and Biborski (2023); Motruk et al. (2023); Wang et al. (2023), but the nontrivial topology invalidates a description as a one-band model. This contrasts MoTe2 bilayers Crépel and Millis (2024b); Qiu et al. (2023) in which more weight lies in the $H_{1,2}$ orbitals, leading to different physical behaviors. In Fig. 1 (b), we display the real-space structure of the three Wannier orbitals. Their exponential localization implies short-ranged hopping parameters as demonstrated in Fig. 1 (a): Disregarding hopping parameters for distances $d\gtrsim 4.5\,|\bm{a}_{M}|$ (with $\bm{a}_{M}$ the moiré lattice vectors) does not change the resulting band structure significantly. A detailed analysis of the energetic and topological properties of the Wannier functions is provided in the SM SM . We carry out the Wannierization for several values of displacement field $E_{z}$ and interpolate the hopping parameters in order to obtain a Hamiltonian for arbitrary $E_{z}$. The resulting density of states is a smooth function of $E_{z}$ [see Fig. 1 (c)], and displays a clear maximum when the Fermi energy meets the VHS of the topmost valence band Schrade and Fu (2021). Along this line, an exceptional point occurs at $E^{c}_{z}\approx 33\,\mathrm{meV}$ (blue circle), where three van-Hove points merge with each other and form a HOVHS Bi and Fu (2021). For $E_{z}<E_{z}^{c}$, the three van-Hove points move towards the $M$-point forming small triangular Fermi pockets around $K^{\nu}$ that are exclusively $T$/$H_{2}$-polarized. In contrast, for $E_{z}>E_{z}^{c}$, the three van-Hove points move towards $\Gamma$, and the Fermi pockets around $K^{\nu}$ disappear. The remaining dispersive part of the Fermi surface is entirely $T$/$H_{2}$-polarized for all values of the displacement field. The multi-orbital Wannier model has the symmetry group $\mathcal{G}_{0}=C_{3z}\otimes U_{\nu}(1)\otimes\mathcal{T}$, where $C_{3z}$ is the point group of the moiré unit cell that contains a three-fold rotation around the $z$-axis, the $U_{\nu}(1)$-symmetry acts in spin-valley space and $\mathcal{T}$ denotes time-reversal symmetry. Due to the absence of cubic terms to lowest order in the continuum description Crépel and Millis (2024b), the system we study theoretically further possesses a mirror-symmetry in each valley $\mathcal{M}_{x}$ that is weakly broken in the physical bilayer but is visible in the fermiology of a single spin-valley sector in Fig. 1 (c).
With the tight-binding Wannier model for arbitrary $E_{z}$ as well as interaction parameters from projection of the dual-gated Coulomb interaction onto the Wannier functions (see SM SM ) at hand, we can formulate the interacting low-energy Hamiltonian as
[…]
where $\nu$ denotes the locked spin-valley degree of freedom, $X,X^{\prime}$ orbital, and $\bm{R},\bm{R}^{\prime}$ Bravais lattice vectors. The operator $c_{\bm{R}X}^{\nu\,(\dagger)}$ destroys (creates) an electron with spin/valley $\nu$ in the orbital $X$ on site $\bm{R}$, and $n^{\nu}_{\bm{R}X}=c^{\nu\,\dagger}_{\bm{R}X}c^{\nu{\vphantom{\dagger}}}_{\bm{R}X}$ is the density operator. As the hopping parameters $t_{\bm{R}X,\bm{R}^{\prime}X^{\prime}}^{\nu}$ are valley-diagonal, the interaction part is what couples different spins/valleys. To disentangle the microscopic mechanisms responsible for the formation of the Fermi surface reconstructed AFM phase and superconductivity observed in experimental measurements Xia et al. (2024); Guo et al. (2024), we first consider the dominant on-site interaction associated to the triangular ($U_{T}$) and honeycomb ($U_{H_{1,2}}$) sites and later investigate the influence of long-ranged density-density interaction on the phase diagram of $\theta=5.08^{\circ}$ tWSe2. The values of $U_{T}\approx 74\,\mathrm{meV}$ and $U_{H}\approx 106\,\mathrm{meV}$ place tWSe2 into the weak-to-intermediate coupling regime, where the band width $W\approx 200\,\mathrm{meV}$ of the three-orbital model fulfills $W\gtrsim U_{T/H}$. We note that due to the topological entanglement of the topmost valence bands in tWSe2, the bandwidth of the entire Wannier band manifold (and not only the topmost valence band) should be considered.
Figure 2: FRG phase diagram of $\theta=5.08^{\circ}$ tWSe2 resolving the interplay of IVC-AFM order and chiral $d/p$-wave superconductivity along the gate-tunable van-Hove singularity. (a) Main panel: FRG phase diagram as function of holes per moiré unit cell $n/n_{0}$ and external displacement field $E_{z}$. The phase diagram shows the critical scale $\Lambda_{\mathrm{c}}$ of the leading Fermi surface instability and indicates regimes of spin/charge density wave order (DW, red) and superconductivity (SC, blue). The density of states (DOS) is shown in the background (gray) to pinpoint the position of the van-Hove line. Close to the HOVHS (blue circle), the system features IVC-AFM order, while for lower values of the displacement field chiral $d/p$-wave superconductivity emerges on the van-Hove line. Inset: Zoom-in on the density-displacement field region where IVC-AFM order gives way to superconductivity. The colored arrows indicate different directions when moving along the van-Hove line (green arrows) or at constant displacement field cuts (grey, blue arrows). The critical scale $\Lambda_{\mathrm{c}}$ along these line cuts is shown in the panels on the left, indicating a maximal critical temperature of $\Lambda_{\mathrm{c}}\sim 500\,\mathrm{mK}$ that is reached in the immediate vicinity of the IVC-AFM phase. (b) Momentum structure and (in)commensurability of the IVC-AFM state along the van-Hove line. The subpanels show the absolute value of the particle-hole susceptibility $\chi^{\text{FRG}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda_{\mathrm{c}}$ for the relevant orbital components involving the Wannier orbitals $T/H_{2}$. The leading momentum transfer $\bm{Q}_{C}$ locks to $K^{\prime}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta along the high symmetry path $K^{\prime}$-$M$ when lowering the external displacement field along the van-Hove line (blue square). (c) Amplitude and phase dependence of the pairing instability at $E_{z}=25\,\mathrm{meV}$ (blue circle) decomposed into singlet $\psi_{0}(\bm{k})$ and triplet $d_{z}(\bm{k})$ components. The continuous colorbar indicates the relative phase, whereas the amplitude is encoded by the opacity. The SC order parameter is two-fold degenerate and likely minimizes its free energy by the chiral superposition $d+id/p+ip$.
II.2 Functional renormalization group approach
In the weak-to-moderate coupling regime, the functional renormalization group (FRG) Salmhofer and Honerkamp (2001); Metzner et al. (2012); Platt et al. (2013); Dupuis et al. (2021); Beyer et al. (2022) represents a well-established method to predict particle-particle and particle-hole instabilities in an unbiased manner. FRG smoothly connects the non-interacting action to the interacting one by virtue of a renormalization group (RG) scale $\Lambda=\infty\to 0$. The introduction of the RG scale leads to an infinite hierarchy of coupled differential equations for the vertex functions $\Gamma^{(2n)}$. In order to numerically solve the coupled nonlinear functional differential equations, we must truncate the hierarchy and approximate the vertex functions’ dependencies on frequency and momentum. To account for the topology of the multi-orbital Wannier setup for the moiré valence bands, we here employ the static four-point truncated unity FRG (TUFRG) approximation Husemann and Salmhofer (2009); Lichtenstein et al. (2017); Profe and Kennes (2022) that allows to keep track of the entire orbital/band and (incommensurate) momentum dependence of the model at the expense of neglecting the frequency dependence of the vertex functions. Most importantly, this allows to treat all couplings on equal footing and (at one loop order) accounts for mutual feedback between the different interaction channels that are renormalized during the flow. The vertex functions are parametrized in a hybrid real-momentum space representation Profe and Kennes (2022); Profe et al. (2024) using a $30\times 30$ momentum mesh for the bosonic momentum variable in each two-particle reducible interaction channel—hence allowing to resolve (in)commensurate particle-particle and particle-hole instabilities—whereas the fermionic momenta are expanded in a real-space form factor basis that is truncated after a distance of 2$|\bm{a}_{M}|$. The propagator is resolved on a refined $900\times 900$ momentum mesh to increase the energy resolution and to sufficiently resolve the fermiology of the system at low scales, see SM SM for details. In static four-point (TU)FRG, the onset of an ordered phase is signaled by the divergence of the two-particle vertex $\Gamma^{(4)}$ at a particular (critical) scale $\Lambda_{\mathrm{c}}$. The most divergent eigenvalue (and the accompanying eigenvector) of $\Gamma^{(4)}$ then provides insights into the structure of the ordered phase—whether it is a particle-particle or a particle-hole instability as well as the momentum, spin/valley and orbital structure of the order parameter. Since we choose the sharp frequency cutoff as regulator (as implemented in the divERGe library Profe et al. (2024)), the critical scale $\Lambda_{c}$ serves as a proxy for the critical temperature of the transition (further details on FRG are given in the SM SM ).
III FRG phase diagram
We present the FRG phase diagram of $5.08^{\circ}$ tWSe2 in Fig. 2 (a), where we plot the critical scale $\Lambda_{\mathrm{c}}$ of emergent particle-particle (blue) and particle-hole (red) instabilities as a function of hole filling per moiré unit cell $n/n_{0}$ and transverse electrical field $E_{z}$. The density of states (DOS) is shown as a reference to pinpoint the position of the van-Hove line (gray). At the position of the HOVHS, the system promotes an inter-valley coherent (IVC) anti-ferromagnetic spin-density wave with order parameter
[…]
where $\vec{\sigma}$ acts in the space of the locked spin-valley degree of freedom and $\vartheta_{X}$ measures the relative orientation and weight of the $120^{\circ}$ in-plane spin orientation on the orbitals $T/H_{2}$ that is otherwise arbitrary due to the remaining $U_{\nu}(1)$ symmetry. The spin density wave breaks translational invariance due to the finite transfer momentum $\bm{Q}_{C}=K^{\nu}$ that connects the HOVHSs in different spin-valley sectors and transforms in the trivial $A$ irreducible representation (IR) of the $C_{3z}$ little group at $K^{\nu}$. Consistent with the orbital polarization of the Fermi pockets at the HOVHS, the order parameter has dominant weight on the $T/H_{2}$-orbitals as imprinted in the transverse particle-hole susceptibility $\chi_{\uparrow\downarrow\uparrow\downarrow}^{\text{FRG}}(\bm{q})$ (see SM SM for details) shown in Fig. 2 (b). Decreasing the value of the displacement field $E_{z}$ along the van-Hove line shifts the leading momentum transfer $\bm{Q}_{C}$ towards the $M$-point, causing regions of incommensurate IVC-AFM order as demonstrated in Fig. 2 (b).
The absence of a perfect scattering vector that connects the van-Hove points in different spin-valley sectors and the reduction of the DOS weakens particle-hole instabilities for $E_{z}<E_{z}^{c}$ and instead promotes the effect of particle-particle fluctuations leading to the formation of superconductivity along the van-Hove line, see blue square in Fig. 2 (a). The formation of superconductivity along the van-Hove line is hence facilitated by Fermi surface warping when detuning the system from the HOVHS, reminiscent of pairing in the $t$-$t^{\prime}$ square lattice Hubbard model Honerkamp (2003) as will be discussed in Section IV.1 in more detail.
The Ising-like spin-orbit coupling in tWSe2 couples singlet- and triplet components of the superconducting order parameter
$\hat{\Delta}^{\text{SC}}(\bm{k})=\left[\hat{\psi}_{0}(\bm{k})\sigma_{0}+\vec{\hat{d}}(\bm{k})\cdot\vec{\sigma}\right]i\sigma_{y}\,,$ (3)
where the hat-symbol encodes the orbital and spin-valley degree of freedom. We find that the leading superconducting order parameter is an inter-valley symmetric combination ($S^{z}=0$) of singlet and triplet components $[\psi_{0}(\bm{k}),d_{z}(\bm{k})]$ that are allowed to admix due to the preserved $U_{\nu}(1)$ symmetry Schrade and Fu (2021). The Fermi surface projected amplitude and phase of the respective components is shown in the right panel of Fig. 2 (c). The leading order parameter transforms in the two-dimensional $E$ irreducible representation of the Hamiltonian’s point group $\mathcal{G}_{0}$. In a mean-field picture, the two degenerate order parameters are in general superimposed and the free energy is presumably minimized by the chiral, time-reversal breaking combination $d\pm id/p\pm ip$. The latter gaps the Fermi surface and hence minimizes the energy of the superconducting condensate Honerkamp (2003); Nandkishore et al. (2012); Black-Schaffer and Honerkamp (2014); Classen et al. (2019); Scherer et al. (2022). The chiral superconducting state belongs to the A Cartan (Altland-Zirnbauer) class for $S_{z}$ preserving Bogoliubov de-Gennes Hamiltonians Schnyder et al. (2008) possessing a $\mathbb{Z}$ topological invariant that manifests in a non-vanishing Chern number $C=\pm 2$ and edge modes that are described by Dirac-like complex fermionic quasiparticle.
Besides the microscopic characterization of the emergent phases, a central question of recent experimental measurements concerns their interplay along the displacement field tunable van-Hove line that is captured by the unbiased FRG analysis. To this end, the inset of Fig. 2 (a) details the transition region of the phase diagram from the intervalley-coherent (IVC) antiferromagnet (AFM) to the chiral $d/p$-wave superconductor. We further provide line cuts of the phase diagram along the van-Hove line (green dash) and for cuts with constant displacement field (blue/gray dash) in Fig. 2 (a). We find that the highest transition temperature of the superconducting state $\Lambda_{c}\sim 500\,\mathrm{mK}$ is reached in the immediate vicinity to the (incommensurate) IVC-AFM phase along the van-Hove line. The superconducting region flanks the IVC-AFM state and we observe an asymmetry of spin-ordered regimes with respect to the van-Hove line, i.e. instabilities occur at larger critical scales $\Lambda_{c}$ for $n<n_{\mathrm{VHS}}$ and $E_{z}<E_{z}^{c}$ in agreement with experimental observations Xia et al. (2024); Guo et al. (2024). The following two Sections are devoted to (i) unraveling the role of IVC-AFM spin-fluctuations in the formation of chiral $d/p$-wave superconductivity and to (ii) tying the observed asymmetry of particle-hole instabilities to the fermiology of the flat bands of tWSe2.
Figure 3: IVC-AFM spin fluctuation mechanism to unconventional superconductivity in tWSe2 by commensurate to incommensurate transition. (a) Absolute value of the RPA-renormalized susceptibility $\chi^{\text{RPA}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at the critical interaction strength $\alpha_{c}U$ involving the dominant Wannier orbitals $T/H_{2}$. The leading momentum transfer $\bm{Q}_{C}$ locks to $K^{\prime}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta along the high symmetry path $K^{\prime}$-$M$ when lowering the external displacement field along the van-Hove line (blue square). (b) Fluctuations around the (incommensurate) IVC-AFM order give rise to chiral $d/p$-wave superconductivity comprising an equal mixture of spin-singlet $\psi_{0}^{\text{RPA}}(\bm{k})$ and spin-triplet $d_{z}^{\text{RPA}}(\bm{k})$ weight in the order parameter. For $E_{z}\sim 25\,\mathrm{meV}$ (blue square), the phase of the resulting order parameter is indicated by the circular colorbar and the amplitude (projected on the Fermi surface) is encoded by the opacity. The superconducting order parameter only has notable weight on the triangular pockets around $K^{\nu}$, whereas the superconducting amplitude vanishes on the larger Fermi arcs connecting neighboring $K^{\nu}$ points.
IV Mechanism to superconductivity
IV.1 Spin-fluctuation exchange as pairing mechanism in tWSe2
To unravel the microscopic mechanism that supports superconductivity along the van-Hove line, we switch off inter-channel coupling of the full FRG flow equations and thereby arrive at a single-channel RPA resummation of the interaction in the crossed particle-hole channel
$C^{\mathrm{RPA}}_{U}(\bm{q})=\hat{U}\hat{\chi}^{\mathrm{RPA}}(\bm{q})\hat{U}=\hat{U}\big{[}\mathbb{1}-\hat{U}\hat{\chi}^{0}(\bm{q})\big{]}^{-1}\,.$ (4)
This procedure segregates spin fluctuations from other possible fluctuations captured by the FRG. The resulting RPA-renormalized particle-hole susceptibility $\hat{\chi}^{\text{RPA}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda=0.1\,\mathrm{meV}$ is shown in Fig. 3 (a). As expected from the Fermi surface sheets in the different spin-valley sectors, c.f. Fig. 1 (c), the dominant ordering vector $\bm{Q}_{C}$ locks to $K^{\nu}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta on the $K^{\nu}$-$M$ line when lowering the external displacement field $E_{z}$. In this regime, the Fermi surface consists of triangular pockets around $K^{\nu}$ featuring three inequivalent van-Hove points associated to the three neighboring $M$-points, see Fig. 1 (c). Scattering between the van-Hove points enhances the particle-hole susceptibility leading to a pronounced commensurate-to-incommensurate transition of the IVC-AFM state along the van-Hove line, which is in good agreement with the interacting particle-hole susceptibility obtained from the full-fledged FRG treatment shown in Fig. 2 (b). To address pairing instabilities originating from fluctuations around the (incommensurate) IVC-AFM order that are present in the paramagnetic normal-state, we construct an effective pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ from the RPA renormalized interaction in the exchange channel Takimoto et al. (2004); Fischer et al. (2021) and analyze the symmetry of the leading superconducting order parameter by solving a linearized gap equation in orbital space, see SM SM for technical details. As shown in Fig. 3 (b), we find a superconducting state with mixed-parity that transforms in the two-dimensional $E$ irreducible representation ($d/p$-wave) in leading order for all values of the displacement field, exemplified for $E_{z}=25\,\mathrm{meV}$ (blue square). The superconducting order parameter is enhanced on the triangular Fermi pockets around $K^{\nu}$ that are exclusively $T/H_{2}$ polarized. Meanwhile, the amplitude of the SC order parameter is significantly suppressed if pairing involves Bloch states associated to the Wannier orbital $H_{1}$. This manifests in a vanishing SC amplitude on the larger Fermi surface arcs that connect neighboring $K^{\nu}$ points, c.f. Fig. 1 (c). Therefore, we expect that the superconducting state is effectively gapless over large parts of the Fermi surface leading to nodal signatures in spectroscopy measurements. The emergence of superconductivity from spin-fluctuation exchange can be understood by the real-space structure of the effective RPA pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ that features attractive components on bonds connecting neighboring $(T,H_{2})$ orbitals such to overcome the initially repulsive Coulomb interaction. As the phenomenology of the IVC-AFM state and the symmetry of the superconducting order parameter accurately agrees with the FRG results, we argue that spin fluctuations represent the pairing glue for electron-mediated superconductivity observed in tWSe2 along the van-Hove line.
Figure 4: Asymmetry of particle-hole instabilities due to Fermi surface broadening induced nesting. (a) Critical (relative) interaction scale $\alpha_{c}$ of the RPA renormalized interaction for different scales $\Lambda$ along the van-Hove line. The leading momentum transfer $\bm{Q}_{C}$ along the high symmetry path $K^{\nu}$-$M$ is encoded by the colorbar shared among panels (a,b). The minimal value of $\alpha_{c}$ for each curve is indicated by the gray dashed line showing a clear asymmetry of particle-hole instabilities with respect to the position of the HOVHS (blue circle). The lower panel shows the ($\Lambda$-broadened) Fermi surface sheets of the spin-valley polarized sub-sectors, i.e., all eigenenergies that are within the energy shell $[E_{F}-\Lambda,E_{F}+\Lambda]$ around the Fermi energy $E_{F}$. (b) Same information as in panel (a), taken for the constant displacement field cut $E_{z}=29\,\mathrm{meV}$ as indicated by the grey arrows in Fig. 2 (a). The grey solid line indicates the DOS as the density $n/n_{0}$ is varied across the van-Hove line. For low scales, the minimal value of $\alpha_{c}$ (grey dashed line) is encountered on the van-Hove line (green arrow) and the leading momentum transfer $\bm{Q}_{c}$ is displaced from $K^{\nu}$. At larger scales, $\alpha^{\text{min}}_{c}$ is shifted to larger hole fillings and $\bm{Q}_{c}$ locks back to $K^{\nu}$.
IV.2 Asymmetry of particle-hole instabilities by Fermi surface broadening induced nesting
An interesting observation of recent experiments Xia et al. (2024); Guo et al. (2024) concerns the apparent asymmetry of instabilities with respect to the displacement field tunable VHS that also manifests in the FRG phase diagram, i.e. (i) the maximal transition temperature of the IVC-AFM phase is encountered for values of the displacement field below the position of the HOVHS and (ii) instabilities are enhanced for $n<n_{\mathrm{VHS}}$, but diminish for $n>n_{\mathrm{VHS}}$. Here, we demonstrate that in the weak-to-moderate interacting regime, this asymmetry is related to scale-dependent nesting properties of the Fermi surface Mai et al. (2022); Šimkovic IV et al. (2022) that lead to commensurability locking of particle-hole instabilities away from the van-Hove line. To analyze the aforementioned asymmetric behavior, we again concentrate on a single-channel RPA resummation in the crossed particle-hole channel and calculate the critical (relative) interaction strength $\alpha_{c}$ defined by the divergence of the RPA renormalized interaction $C^{\mathrm{RPA}}_{U}(\bm{q})$ [cf. Eq. 4], i.e., $\det[\mathbb{1}-\alpha_{c}\hat{U}\hat{\chi}^{0}(\bm{q})]=0$. Figure 4 (a) displays the critical value $\alpha_{c}$ and the leading momentum transfer $\bm{Q}_{C}$ of the (incommensurate) IVC-AFM state along the van-Hove line for different RG scales $\Lambda$. The latter enters the RPA theory as temperature in the Fermi-Dirac statistics and hence determines the smearing of states in the vicinity of the Fermi energy, see SM SM for technical details. At low scale $\Lambda=0.1\,\mathrm{meV}$, the minimal critical interaction $\alpha_{c}^{\text{min}}$ (gray dashed line) is found at the HOVHS (blue circle), where the leading momentum transfer of the associated IVC-AFM state locks to $\bm{Q}_{C}=K^{\nu}$. This finding is consistent with the enhanced DOS at the HOVHS and the fact that the vector $\bm{q}=K^{\nu}$ perfectly connects Fermi surface sheets (FS) of opposite spin-valley polarization. To this end, we show the $\Lambda$-broadened FS, i.e., all eigenenergies that reside within the energy shell $[E_{F}-\Lambda,E_{F}+\Lambda]$ around the Fermi energy $E_{F}$, in the lower panel of Fig. 4 (a). Lowering the displacement field continuously displaces $\bm{Q}_{C}$ towards the $M$-point and increases the critical interaction strength $\alpha_{c}$ as scattering between the triangular FS pockets around $K^{\nu}$ is suppressed due to the relative orientation mismatch of the van-Hove points in different spin-valley sectors. At larger scales $\Lambda=\{0.5,1.0\}\,\mathrm{meV}$, however, the suppression of particle-hole instabilities for $E_{z}<E_{z}^{c}$ and the associated shift of $\bm{Q}_{C}$ towards the $M$-point are counteracted by the induced smearing of the triangular FS pockets formed by the three van-Hove points. Lowering the displacement field along the van-Hove line (left panel) increases the size of the FS pockets around $K^{\nu}$ such that scattering with vector $\bm{Q}_{C}=K^{\nu}$ between FS sheets of opposite spin-valley polarization is (re)enhanced at larger scales $\Lambda$. As a consequence of the higher scale $\Lambda$, the minimal critical value $\alpha_{c}^{\mathrm{min}}$ moves to lower values of the displacement field along the van-Hove line. This causes an asymmetry of interaction-induced particle-hole instabilities with respect to the HOVHS.
A similar asymmetry is encountered when varying the hole density $n/n_{0}$ for constant values of the displacement field $E_{z}<E_{z}^{c}$ as demonstrated in Fig. 4 (b). To show this, we focus on $E_{z}=29\,\mathrm{meV}$ (gray dash), where the asymmetry is maximal in the FRG phase diagram as demonstrated in Fig. 2 (a). In this regime, the Fermi surface at the VHS ($n_{\mathrm{VHS}}\sim-1.3$, green arrow) consists of three van-Hove points that are displaced from $K^{\nu}$ such that the leading momentum transfer $\bm{Q}_{C}$ is shifted towards the $M$-point. Decreasing the hole density lifts the van-Hove points and increases the size of the (open) FS pockets, similar to the observed phenomenology when varying the displacement field along the van-Hove line. For larger scales (temperatures), the broadening of the FS pockets (re)enhances scattering processes with momentum transfer $\bm{Q}_{C}=K^{\nu}$ and the minimal critical interaction strength $\alpha_{c}^{\text{min}}$ is moved to hole densities $n<n_{\mathrm{VHS}}$.
In conclusion, we argue that the minimal critical interaction $\alpha_{c}^{\text{min}}$ is shifted to lower displacement fields $E_{z}<E_{z}^{c}$ and lower densities $n<n_{\mathrm{VHS}}$ with respect to the position of the HOVHS and the displacement field tunable van-Hove line when particle-hole instabilities are encountered at non-zero temperature, or higher scales $\Lambda$. As the critical scale $\Lambda_{c}$ of particle-hole instabilities encountered in the FRG flow depends primarily on the strength of the Hubbard interactions $U_{X}$ in Eq. 1, it is an indicator for how strongly coupled the system under scrutiny is. Therefore, it is expected that the observed asymmetry is enhanced when increasing the strength of the interactions relative to the bandwidth. This can be achieved by tuning the twist angle (see Section VI) or altering the dielectric environment, i.e., changing the distance to metallic gates. We further note that the scale-dependent nesting properties of tWSe2 are tied to microscopic details of the fermiology of the system, particularly the formation of the triangular FS pockets around $K^{\nu}$ and the flat regions in momentum space originating from the HOVHS. This asymmetric occurrence of particle-hole instabilities and the associated commensurability locking may hence not be unique to tWSe2, but appear generically in systems with tunable FS pockets and HOVHS.
Figure 5: Influence of long-ranged density-density interactions on the low-energy phase diagram of $\theta=5.08^{\circ}$ tWSe2. (a) FRG phase diagram as function of holes per moiré unit cell $n/n_{0}$ and external displacement field $E_{z}$. The on-site IVC-AFM order parameter mixes with a spin-bond order (SBO) parameter between the $T$/$H_{2}$ orbitals that transform in the $A$-IR of $C_{3z}$. This raises the critical scale $\Lambda_{\mathrm{c}}$ of the density wave (DW) instabilities in the phase diagram and enhances the critical temperature of the superconducting phase. (b) Real-space structure of the particle-hole instabilities. The (magnetic) Wigner-Seitz cell is indicated by the (green) gray dashed line.
V Influence of long-ranged interactions: Formation of spin-bond order
Besides the dominant Hubbard interaction terms acting on the triangular ($T$) and honeycomb sites ($H_{1,2}$) of the effective Wannier model defined in Eq. 1, the ratio between the Wannier function spread $\Omega\sim|\bm{a}_{M}|=3.74\,\mathrm{nm}$ and the distance to the metallic gates $\xi\sim 10\,\mathrm{nm}$ on either side of the sample suggests that longer-ranged density-density interactions impact the low-energy physics of $\theta=5.08^{\circ}$ tWSe2. Meanwhile, exchange and pair-hopping contributions between neighboring Wannier orbitals are negligible as expected from the exponential localization of the Wannier orbitals, see SM SM . Therefore, we next complement the Hubbard Hamiltonian Eq. 1 by density-density terms that are obtained when projecting the dual-gated Coulomb interaction to the Wannier basis:
$H^{V}=H+\sum_{\nu,\mu}\!\!\!\sum_{\begin{subarray}{c}X,X^{\prime}\in\\ \{T,H_{1},H_{2}\}\end{subarray}}\!\!\sum_{\bm{R},\bm{R}^{\prime}}V_{XX^{\prime}}^{\bm{R}\bm{R}^{\prime}}n^{\nu}_{\bm{R}X}n^{\mu}_{\bm{R}^{\prime}X^{\prime}}\,.$ (5)
In the spirit of an ab-initio characterization of correlated phases in tWSe2, we consider the full real-space dependence of $V_{XX^{\prime}}^{\bm{R}\bm{R}^{\prime}}$ in this Section. The largest of these density-density terms are the nearest-neighbor honeycomb-triangular (honeycomb-honeycomb) density-density interactions $V_{TH}\approx 33\,\mathrm{meV}$ ($V_{HH}\approx 28\,\mathrm{meV}$). We demonstrate in the SM SM that truncating Eq. 5 after these terms already qualitatively accounts for the effects of longer-range interactions observed below.
The corresponding FRG phase diagram for $\theta=5.08^{\circ}$ tWSe2 is shown in Fig. 5 (a). While the qualitative features of the phase diagram resemble those of the pristine Hubbard model Eq. 1, the phase boundaries and critical scales of the emergent particle-hole (particle-particle) instabilities are altered. The observed quantitative changes arise because the inter-site interaction causes non-local particle-hole pairs to get admixed to the local IVC-AFM order parameter. The nearest-neighbor repulsion between triangular ($T$) and honeycomb ($H_{2}$) orbitals supports the formation of a spin/valley-bond order (SBO)
$\vec{\Delta}^{\mathrm{SBO}}_{\bm{R},\bm{R}^{\prime}}\propto e^{i\bm{Q}_{C}\cdot\bm{R}}\braket{c^{\nu\dagger}_{\bm{R},T}\vec{\sigma}^{\nu\nu^{\prime}}c^{\nu^{\prime}{\vphantom{\dagger}}{}}_{\bm{R}^{\prime},H_{2}}}+\mathrm{h.c.}$ (6)
The SBO transforms within the trivial $A$-irreducible representation (IR) of the little group $C_{3z}$ at $K^{\nu}$ and is therefore symmetry-allowed to mix with the IVC-AFM state $\Delta^{\mathrm{AFM}}$ driven by the on-site Hubbard repulsion. In Fig. 5 (b), we present a sketch of the magnetization patterns in the IVC-AFM+SBO phase comprising an on-site $120^{\circ}$ spin configuration as well as the bond order component between $T$-$H_{2}$. Note that due to the $p_{\nu}$-orbital character of the $H_{2}$-orbital, the bonds feature a relative phase of $\omega=e^{2\pi i/3}$ between $C_{3z}$-related $H_{2}$-orbitals. The gain in free energy arises because the projection of the bare nearest-neighbor repulsion $V_{TH(HH)}$ to the crossed particle-hole channel lowers the energy of states with an SBO component of the order:
$\Delta^{\text{SBO}*}\circ C\big{[}\Gamma^{(4)}_{\Lambda=\infty}\big{]}\circ\Delta^{\text{SBO}}>0\,,$ (7)
see SM SM for details. The non-vanishing SBO components in the order parameter hence promote the formation of particle-hole pairs at larger critical scales $\Lambda_{c}$ as demonstrated in Fig. 5 (a) [compare with Fig. 2 (a)]. We therefore argue that the particle-hole instability experiences an effective anti-screening by virtue of long-ranged density-density type interactions. The coupling to the on-site IVC-AFM order that drives the chiral $p/d$-wave pairing instability causes the maximal critical temperature to rise to $\Lambda_{c}\sim 1\,\mathrm{K}$. Meanwhile, the size of the SC region in the phase diagram is decreased due to the competition with the IVC-AFM state.
Figure 6: Evolution of IVC-AFM domains and SC order in the low-energy phase diagram of tWSe2 in the intermediate twist angle regime $\theta=4^{\circ}\dots 5^{\circ}$. (a) FRG phase diagram at $\theta=4.5^{\circ}$ tWSe2 considering Hubbard-type interactions on the triangular ($T$) and honeycomb ($H_{1,2}$) orbitals. The blue circle indicates the position of the HOVHS. Regions of IVC-AFM order evolve to smaller absolute values of the displacement field $E_{z}$ and larger critical scales signaling the onset of correlated states at higher critical temperatures $\Lambda_{\mathrm{c}}\sim T_{\mathrm{c}}$. The maximal critical scale $\Lambda^{\text{max}}_{\mathrm{c}}$ features a pronounced asymmetry with respect to the HOVHS (blue circle) and the tunable van-Hove line (grey). (b) FRG phase diagram at $\theta=4.1^{\circ}$. Reaching the moderate-to-strongly interacting regime, the critical scales of IVC-AFM and chiral $d/p$-wave superconductivity are further enhanced. As a consequence, regions of SC order in the $E_{z}-n/n_{0}$ phase diagram are shrunk in size in favor of IVC-AFM ordered regions that move towards $n/n_{0}\sim 1$.
VI Evolution of the SC phase in the intermediate twist angle regime
Finally, we map out phase diagrams in the intermediate twist angle regime $\theta=\{4.1^{\circ},4.5^{\circ}\}$ as summarized in Fig. 6 (a,b). Decreasing the twist angle quenches the kinetic energy scales such that the moiré bands react more sensitively to the external displacement field. As a result, the position of the HOVHS is moved to smaller values of $E_{z}$, see blue circle in Fig. 5 (c,d) and SM SM . As the ratio between the Wannier function spread and the gate screening distance $\Omega/\xi$ decreases with the twist angle, long-ranged density-density interaction between the Wannier orbitals are successively suppressed and we therefore resort to the Hubbard Hamiltonian defined in Eq. (1). The values of the Hubbard interactions for different twist angles are listed in the SM SM .
Apart from the enhanced impact of the displacement field on the single-particle bandstructure, decreasing the twist angle increases the ratio $U/W$ and places the system in a moderate interacting regime. As a consequence, the emergent IVC-AFM state and the chiral $d/p$-wave SC state manifest at larger scales $\Lambda_{c}$ indicating a higher critical temperature of respective phases. Consistent with the analysis brought forward in Section IV.2, this enhances the asymmetry of the IVC-AFM with respect to the van-Hove line as demonstrated in Fig. 6 (a,b). At $\theta=4.5^{\circ}$, the maximal critical scale $\Lambda_{c}$ of the IVC-AFM phase is found around $E_{z}\sim 15\,\mathrm{meV}$ and is clearly displaced from the van-Hove line, where instead regions of chiral $d/p$-wave SC prevail. At $\theta=4.1^{\circ}$ the IVC-AFM state is displaced to the vicinity of $n/n_{0}\sim 1$ and the superconducting region is shrunk in size; a trend that is observed in small twist angle tWSe2 in Ref. Xia et al. (2024). In particular, superconductivity does not extend to $E_{z}=0\,\mathrm{meV}$, but is tied to an intermediate displacement field regime $5\,\mathrm{meV}<E_{z}<10\,\mathrm{meV}$.
VII Conclusion and outlook
In this work, we provide a complete and unbiased characterization of correlated states in $\theta=4^{\circ}\dots 5^{\circ}$ tWSe2 from first principles by studying the impact of gate-screened Coulomb interactions with state-of-the-art functional renormalization group techniques. Our analysis reveals that inter-valley coherent anti-ferromagnetic order (IVC-AFM) and chiral, mixed-parity $d/p$-wave superconductivity collaborate along the gate-tunable van-Hove line. The superconducting pairing glue for the unconventional SC state is provided by spin-fluctuations arising in regions of incommensurate IVC-AFM order below the HOVHS, where the Fermi surface pockets around the valleys $K^{\nu}$ are broadened and therefore weaken scattering between opposite spin-valley sectors. We attribute the asymmetry of particle-hole instabilities observed in experiment Xia et al. (2024); Guo et al. (2024) to the formation of hole pockets around $K^{\nu}$ that are broadened at larger scales (temperature) such that the dominant ordering vector locks back to $K^{\nu}$ at positions that are offset from the van-Hove line  [^1] We foresee this scale-dependent nesting as a general feature of weak-to-moderately interacting fermions with HOVHS.. As the hole pockets are equally polarized on the $T/H_{2}$-orbital, this stresses the importance of describing the low-energy physics of tWSe2 in the multi-orbital setup presented in this manuscript. One aspect of the experiment of Guo et. al. Guo et al. (2024) that is however not clearly reproduced by our calculations is the non-monotonic variation of antiferromagnetism with carrier concentration at fixed displacement field. At displacement fields slightly greater than those where superconductivity is observed, experiment finds evidence for antiferromagnetism at carrier concentrations slightly lower than and slightly higher than the van-Hove concentrations, but not when the carrier density is tuned to the van-Hove concentration. Understanding this is behavior is an important topic for future research. Further, we demonstrate that long-ranged density-density interactions do not alter the phase diagram of tWSe2 qualitatively, but change the phase boundaries and critical temperature due to the coupling of the IVC-AFM order parameter to a spin/valley bond order (SBO) that acts as an interaction-induced spin-orbit coupling term. We therefore argue that the presence of longer-ranged density-density interactions will enhance the critical temperature $T_{c}$ of the SC phase, however, decrease its overall size in the phase diagram due to the competition with the IVC-AFM state.
Finally, we show that as the twist angle is decreased to $\theta=4.1^{\circ}$ the regions of SC order are displaced to smaller values of the external displacement field ($5\,\mathrm{meV}<E_{z}<10\,\mathrm{meV}$) and to hole densities closer to $n/n_{0}\sim 1$ in good agreement with experimental measurements. It is interesting to consider these results in light of the finding of Xia et. al. Xia et al. (2024) that in a $\theta=3.65^{\circ}$ angle device, a superconducting region exists very near $n/n_{0}=1$ and vanishing displacement field, giving way to an insulating phase as the displacement field is increased. Whether this superconductor-insulator transition is a consequence of strong coupling Mott/spin liquid physics or can be understood within the weak-to-moderate coupling physics captured by the FRG approach used here is an important open question.
The validity of the static four-point FRG employed in this work is bound to the weak to moderate interacting regime $U/W<1$ that is relevant for the twist angles considered within this work. By further decreasing the twist angle, we expect that the system is eventually tuned into a strongly interacting regime. It is thus a highly anticipated avenue of future research to treat this regime with state-of-the art methods such as, e.g. DMFT or slave Bosons to learn about the properties of strongly correlated and ordered states in tWSe2.
Acknowledgements.
This work was supported by the Excellence Initiative of the German federal and state governments, the Ministry of Innovation of North Rhine-Westphalia and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). AF and DMK acknowledge funding by the DFG within the Priority Program SPP 2244 “2DMP” – 443274199. LK acknowledges support from the DFG through Project-ID 258499086 – SFB 1170 and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat, Project-ID 390858490 – EXC 2147. LK and TOW greatfully acknowledge support from the DFG through FOR 5249 (QUAST, Project No. 449872909) and SPP 2244 (Project No. 422707584). SR is supported by the DFG research unit FOR 5242 (WE 5342/7-1, project No. 449119662). TOW is supported by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the DFG (EXC 2056, Project ID 390715994). AR acknowledges support by the European Research Council (ERC-2015-AdG694097), the Cluster of Excellence ‘Advanced Imaging of Matter’ (AIM), Grupos Consolidados (IT1249-19) and Deutsche Forschungsgemeinschaft (DFG) – SFB-925 – project 170620586. LX acknowledges supported by the National Key Research and Development Program of China (Grant No. 2022YFA1403501), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2022B1515120020), the National Natural Science Foundation of China (Grant No. 62341404), Hangzhou Tsientang Education Foundation and the Max Planck Partner group programme. AJM acknowledges support from Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award (DE-SC0019443). DMK and AR acknowledge support by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. Computations were performed on the HPC system Ada at the Max Planck Computing and Data Facility. The Flatiron Insitute is a division of the Simons Foundation.